OpenStudy (anonymous):

A fire on A 150ft. Tall tower Spots A Fire At a 30degree angle of depression. About how many feet away from this tower was the fire. ( i have 4 multiple choice answers : 86.6ft. , 129.9ft , 259.8ft. & 300ft.)

OpenStudy (dumbcow):



OpenStudy (nurali):

Diagram as a right triangle height = 150ft, d the other leg and a 30 degree angle opposite d. tan30 = d/150 150(0.577) = d d = 86.6 feet

OpenStudy (dumbcow):

No opp angle is 60 tan(60) = d/150 d = 150sqrt(3)

OpenStudy (anonymous):

thank you @Nurali

OpenStudy (nurali):

Anytime.